Skip to Search

diagonal spacer

Systems Staff Meeting Notes
October 24, 2005

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Geoffrey Stewart, Mark Sullivan, Ying Tang
  1. Tom saw the worse movie of the year (so far) over the weekend: "Elizabethtown".
  2. "Doom" also scored 28% on the Rotten Tomato meter.
  3. Break room: refrigerator door not closing due to iceberg growing in the freezer. It is currently being defrosted. Also, the snacks from Friday were left out all weekend. It should be the duty of the person who brings snacks to clean up by 5PM Friday. Geoffrey will check into the possibility of getting a new refrigerator via surplus property. Will's signature will be required.
  4. Town Meeting November 17 - all staff are encouraged to go. There will be employee recognition.
  5. Library Council - discussed how FICA will no longer be paid for OPS employees. UF contributions to be used for savings - 401B account.
  6. Staff position classification review - going over positions to make them fair across the board. Will be at least 6 months before it done. Staff will have the opportunity to contribute.
  7. New procedure for use of public folders - Systems approval now required. Systems will be the one to create the folders. There will be an acceptable use policy.
  8. AD migration is complete. We are in full AD mode - no more mixed mode, no going back. So far, Web, Exchange and Domain Controllers are al upgraded and appear to be fine. Winston notes there was a campus-wide AD meeting scheduled on Friday, but it was cancelled at the last minute.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Two boxes were misdirected on recently. Debra called Brian in Facilities - he says they have a new person delivering and he will rectify.
  1. Public computer configuration - Directors will look to hire temporary IT staff to complete the project.
  1. McAfee - Will to put on the servers and Geoffrey to work on the desktop version.
  2. Greenstone Production server - Geoffrey to work on it. Ditto for BBS.
Software Projects Status
  1. Patron Updates - new address info went out successfully on today's load.
  2. Greenstone - same.
  3. Programmer position - Winston is reviewing resumes.
Web Coordinator
  1. Tom sent out instructions on Friday for Web page uploads to the new server. The config appears to be different on some machines.
  2. Unpublished drafts on the Web site are gone. May be able to find them on the local machine. You can export them to the desktop, kill the site and import them back.
  3. NT8 is gone: unplugged. Almost everything appears to be working. NT12 ended up giving us the most permissions problems.
Security Meeting
  1. Meeting was held.


University of Florida Home Page